OpenStudy (anonymous):

Simplify the expression below. 6n2 - 5n2 + 7n2

Nnesha (nnesha):

first step find GCF(greatest common factor)

OpenStudy (anonymous):

Are those squared?

OpenStudy (anonymous):

yeah

Nnesha (nnesha):

combine like terms

OpenStudy (michele_laino):

hint: \[\large 6{n^2} - 5{n^2} + 7{n^2} = {n^2}\left( {6 - 5 + 7} \right) = ...?\]

OpenStudy (anonymous):

8n^2

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

all variables are same and same exponent so you can combine them \[\huge\rm 6\color{green}{n^2} -5\color{green}{n^2} + 7\color{green}{n^2}\]

OpenStudy (michele_laino):

thanks